# SpotLight: Shadow-Guided Object Relighting via Diffusion

### v2

@@ 5.2 SpotLight specifics
As mentioned in sec. 3, our proposed SpotLight readily applies to existing diffusion-based neural renderers, such as ZeroComp [78] or RGBX [75]. Here, we provide details on ZeroComp, see supp. for RGBX.
We use the checkpoint trained on the Openrooms dataset [37] provided by the authors. The masked shading map is obtained by dividing the background image by the albedo, and by masking out the shading on both the object and shadow regions. Given its superior results, we adopt this method as our default model for 3D object compositing. To extract intrinsics from real images, we use the same conditioned intrinsic estimators as in [78], namely StableNormal [71] for normals, IID [32] for albedo, and ZoeDepth [8] for depth.
@@
| DiLightNet [74] | 24.67 | 0.948 | 0.064 | 0.022 | 0.042 |
| Neural Gaffer [29] | 28.44 | 0.963 | 0.042 | 0.015 | 0.038 |
@@
| IC-Light [76] | 26.87 | 0.959 | 0.054 | 0.019 | 0.040 |
| ZeroComp+SDEdit [44] | 26.00 | 0.938 | 0.053 | 0.025 | 0.079 |
@@ 5.3 Baselines
We compare SpotLight with two types of baselines, that is methods that are conditioned on: the guiding shadow (like SpotLight); or an explicit lighting representation, through a environment map. We now describe per-method details. OpenImageDenoise [1] is also used to denoise the results of all methods to limit the impact of noise generated by diffusion models.
@@ 5.3.1 Shadow-conditioned methods
ZeroComp [78]+SDEdit [44] We create a strong baseline by first running ZeroComp with the shading of the object masked out, and composite the guiding shadow over the background shading intrinsic map. To blend the shadows, we perform a refinement step, inspired by SDEdit [44], by noising this prediction for 50% of the timesteps, and re-running ZeroComp for the remaining timesteps, this time with the shading of the whole image masked out.
IC-Light [76] We evaluate the background-conditioned model of IC-Light, where the background image with the shadow is used for conditioning. We provide a rendering of the object lit by constant ambient lighting as input. No prompts are used.
@@ 5.3.2 Light-conditioned methods
DiLightNet [74] DiLightNet requires a prompt describing the object, which we obtain by feeding the object lit by constant ambient lighting and composited on a black background to BLIP-2 [35] with the prompt “Putting aside the black background, this object is a”. DiLightNet also requires depth and radiance hints: we obtain depth from the 3D model of the object and render the radiance hints using the input environment map as outlined in [74]. The rendered object is then composited with the background.
Neural Gaffer [29] The object lit by constant ambient lighting is composited on a white background before feeding it to the network, as expected by the method. Since Neural Gaffer is trained at a lower resolution (), its result is upsampled to then composited on the full resolution background. We noticed Neural Gaffer tends to generate noticeable white boundary artifacts—we remove them through an erosion operation to avoid overly penalizing it.
| Guiding shadow | IC-Light [76] | DiLightNet [74] | Neural Gaffer [29] | ZeroComp+SDEdit | Ours | Ground truth |
@@ 5.5 User studies
Two user studies were conducted to assess the perceptual realism and controllability of SpotLight compared to baselines. In both cases, the Thurstone Case V Law of Comparative Judgement [61] is used to obtain a scaling -score for each method, where a higher scale indicates a higher human preference. Here, we give a high-level overview of those user studies. Refer to the supplementary material for additionnal details and studies.
@@
Here, we use the “user-controlled” version of the dataset (see sec. 5.1) to evaluate the realism of results obtained by conditioning on a specific lighting direction, even if it is not aligned with the scene lighting conditions. We compare SpotLight against the three baselines with the best PSNR from tab. 1, namely Neural Gaffer [29], IC-Light [76], and ZeroComp [78]+SDEdit [44]. Since no ground truth is available, we design a two-alternative forced choice user study, where results obtained by two methods are shown side by side. The participants are asked to “Click on the image that looks the most realistic, by considering both the appearance of the object and its shadows.” For each pair of methods, a set of 20 random images are chosen (we ensure images where shadows are clearly visible are selected), resulting in 120 pairs of images observed by each participant. We randomly shuffle the 120 pairs for each user, as well as the left-right order. Note that we also add three sentinel images to discard observers that did not understand the task. Examples are shown in fig. 3.
User study I was completed by observers. The results, shown in fig. 4a, show that our method is the favorite, indicating that SpotLight achieves the highest realism, with a statistically significant difference to Neural Gaffer and IC-Light.
@@
User study II was completed by observers. As seen in fig. 4b, SpotLight achieves a statistically significant improvement in performance over ZeroComp+SDEdit, despite the low number of participants, which we attribute to the use of our local guidance strategy.
@@ 5.6 Effect of parameter selection
Our method employs parameters that control the degree of shadow guidance ( in eq. 3) and the shadow blending ( in eq. 2). Tab. 2 shows the impact of parameter selection on quantitative metrics, obtained on the “reference-based” dataset where the ground truth is available (c.f. sec. 5.1). Although using no guidance () results in better quantitative performance, we observe qualitatively in fig. 5 that this significantly reduces the visibility of generated shading on the inserted object, thereby reducing the impact of the desired local lighting control. Employing no shadow blending () results in similar metric values but makes shadows much less visible (refer to the supp. for a qualitative comparison).
@@ 6 Applications
Scribbles as shadows Shadows can be generated in multiple ways: we have demonstrated the use of shadow mapping in 3D and, in the supp., pixel height in 2D [58]. However, one can simply draw the desired shadow! Fig. 6 shows that SpotLight is able to realistically relight the object and refine its shadow even when it is a user-drawn scribble.
Full image relighting As shown in fig. 7, we can extend SpotLight to achieve full-scene relighting. Although our method is not explicitly designed for this task, we find our shadow guidance to be effective in plausibly relighting the background. First, we relight the object using SpotLight with the desired guiding shadow. Next, we apply a version of ZeroComp trained with an inverted shading mask, using only the object shading to relight the background. We repeat this process for different light directions and average the resulting predictions to obtain an average relit background. Finally, we compute a per-pixel highlight map by dividing each background prediction by the average and multiplying it by the original background. This process ensures that background shadows and highlights consistently align with the object’s shadow.
@@ 7 Discussion
| Incorrect user-defined shadow | Correct shadow |
Limitation: physically incorrect shadows Since SpotLight imposes no constraints on the input shadow, the user is free to specify a shadow direction that potentially contradicts the real lighting direction in the background image. In this case, SpotLight will still attempt to generate an image that looks as realistic as possible, given the shadow conditioning, as shown in fig. 8.
We present SpotLight, a method for realistically controlling the local lighting of an object through a coarse shadow, compatible with diffusion-based intrinsic image renderers without requiring any additional training. Our results demonstrate that fine-grained control over local lighting can be achieved while attaining realistic compositions.
While the model can realistically generate relighting results from a shadow and despite showing that this can be done in multiple ways (shadow mapping, 2D generation, hand-drawn scribbles), the process of generating a shadow from scratch could prove challenging to some. A promising future direction is to enable end-to-end shadow control using parametric light models, such as point or sphere lights. Additionally, similar approaches could be explored to enable global lighting adjustments for entire images.
Acknowledgements
This research was supported by NSERC grants RGPIN 2020-04799 and ALLRP 586543-23, Mitacs and Depix. Computing resources were provided by the Digital Research Alliance of Canada. The authors thank Zheng Zeng, Yannick Hold-Geoffroy and Justine Giroux for their help as well as all members of the lab for discussions and proofreading help.

### v3

@@ 5.2 SpotLight specifics
As mentioned in sec. 3, our proposed SpotLight readily applies to existing diffusion-based neural renderers, such as ZeroComp [zhang2025zerocomp] or RGBX [zeng2024rgb]. Here, we provide details on ZeroComp, see supp. for RGBX.
We use the variant trained on Openrooms [li2020openrooms] provided by the authors. The masked shading map is obtained by dividing the background image by the albedo, and by masking out the shading on both the object and shadow regions. Given its superior results, we adopt this method as our default model for 3D object compositing. To extract intrinsics from real images, we use the same conditioned intrinsic estimators as in [zhang2025zerocomp], namely StableNormal [ye2024stablenormal] for normals, IID [kocsis2023iid] for albedo, and ZoeDepth [bhat2023zoedepth] for depth.
@@
| DiLightNet [zeng2024dilightnet] | 24.67 | 0.948 | 0.064 | 0.022 | 0.042 |
| Neural Gaffer [jin2024neural_gaffer] | 28.44 | 0.963 | 0.042 | 0.015 | 0.038 |
@@
| IC-Light [zhang2025iclight] | 26.87 | 0.959 | 0.054 | 0.019 | 0.040 |
| ZeroComp+SDEdit [meng2022sdedit] | 26.00 | 0.938 | 0.053 | 0.025 | 0.079 |
@@ 5.3 Baselines
We compare SpotLight with two types of baselines, that is methods that are conditioned on: the guiding shadow (like SpotLight); or an explicit lighting representation, through a environment map. We now describe per-method details. OpenImageDenoise [OpenImageDenoise] is also used to denoise the results of all methods to limit the impact of rendering noise present in the synthetic training data of some models.
@@ 5.3.1 Shadow-conditioned methods
ZeroComp [zhang2025zerocomp]+SDEdit [meng2022sdedit]. We create a strong baseline by first running ZeroComp with the shading of the object masked out, and composite the guiding shadow over the background shading intrinsic map. To blend the shadows, we perform a refinement step, inspired by SDEdit [meng2022sdedit], by noising this prediction for 50% of the timesteps, and re-running ZeroComp for the remaining timesteps, this time with the shading of the whole image masked out.
IC-Light [zhang2025iclight]. We evaluate the background-conditioned model of IC-Light, where the background image with the shadow is used for conditioning, with the object lit by a constant ambient lighting as input. No prompts are used.
@@ 5.3.2 Light-conditioned methods
DiLightNet [zeng2024dilightnet] requires a prompt describing the object, which we obtain by feeding the object lit by constant ambient lighting and composited on a black background to BLIP-2 [li2023blip2] with the prompt “Putting aside the black background, this object is a”. DiLightNet also requires depth and radiance hints: we obtain depth from the 3D model of the object and render the radiance hints using the input environment map as outlined in [zeng2024dilightnet]. The rendered object is then composited with the background.
Neural Gaffer [jin2024neural_gaffer]. The object lit by constant ambient lighting is composited on a white background before feeding it to the network, as expected by the method. Since Neural Gaffer is trained at a lower resolution (), its result is upsampled to then composited on the full resolution background. We noticed Neural Gaffer tends to generate noticeable white boundary artifacts—we remove them through an erosion operation to avoid overly penalizing it.
@@ 5.5 User studies
Two user studies were conducted to assess the perceptual realism and controllability of SpotLight compared to baselines. In both cases, the Thurstone Case V Law of Comparative Judgement [thurstone2017] is used to obtain a -score for each method, where a higher value indicates higher human preference. Here, we give an overview of those user studies. Refer to the supp. for additional details and studies.
@@
Here, we use the “user-controlled” version of the dataset (see sec. 5.1) to evaluate the realism of results obtained by conditioning on a specific lighting direction, even if it is not aligned with the scene lighting conditions. We compare SpotLight against the three baselines with the best PSNR from tab. 1, namely Neural Gaffer [jin2024neural_gaffer], IC-Light [zhang2025iclight], and ZeroComp [zhang2025zerocomp]+SDEdit [meng2022sdedit]. Since no ground truth is available, we design a two-alternative forced choice user study, where results obtained by two methods are shown side by side. The participants are asked to “Click on the image that looks the most realistic, by considering both the appearance of the object and its shadows.” For each pair of methods, a set of 20 random images are chosen (we ensure images where shadows are clearly visible are selected), resulting in 120 pairs of images observed by each participant. We randomly shuffle the 120 pairs for each user, and the left-right order. Three sentinel images were added to discard observers that did not understand the task. Examples are shown in fig. 3.
User study I was completed by observers. Fig. 4a shows that our method is the favorite, indicating that SpotLight achieves the highest realism, with a statistically significant difference to Neural Gaffer and IC-Light.
@@
User study II was completed by observers. As seen in fig. 4b, SpotLight achieves a statistically significant improvement in performance over ZeroComp+SDEdit, despite the low number of participants, which we attribute to the use of our local guidance strategy. Please see the supp. for more user study results and experimental details.
@@ 5.6 Effect of parameter selection
Our method uses two tunable parameters: the local guidance scale ( in eq. 2) and the shadow blending weight ( in eq. 1). Tab. 2 reports metrics obtained on the “reference-based” dataset where the ground truth is available (c.f. sec. 5.1). Although using no guidance () results in better quantitative performance, we observe qualitatively in fig. 5 that this significantly reduces the visibility of generated shading on the inserted object, thereby reducing the impact of the desired local lighting control. Employing no shadow blending () results in similar metric values but makes shadows much less visible. Please refer to the supp. for a qualitative comparison for the , and the shadow softness (light radius) parameters.
@@ 6 Applications
Scribbles as shadows. Instead of relying on shadow mapping in 3D, one can simply draw the desired shadow! Fig. 6 shows that SpotLight can realistically relight the object and refine its shadow even when it is a user-drawn scribble.
2D objects. SpotLight can be applied to real 2D objects from existing photographs by estimating their intrinsics, as in sec. 5.2, and generating desired shadows with shadow generation methods such as [tasar2024controllable], as illustrated in fig. 7.
| Input object | Relit object on target background | ||
Reflective materials. When using a ZeroComp backbone pre-trained on the InteriorVerse dataset [zhu2022learning], SpotLight naturally generalizes to metallic objects and generates controllable specular reflections on the object, see fig. 8.
Additional relighting results. SpotLight also shows other capabilities: relighting for more diverse objects and outdoor scenes, simulating multiple light sources by combining outputs with different guiding shadows, and full-image relighting. Due to space constraints, the implementation details and results are provided in the supp.
@@ 7 Discussion
| Incorrect user-defined shadow | Correct shadow | ||
Limitation: physically incorrect shadows. Since SpotLight imposes no constraints on the input shadow, users may specify a shadow direction that potentially contradicts the real lighting in the background. In this case, SpotLight still attempts to render a visually plausible composition consistent with the given shadow, as shown in fig. 9.
We present SpotLight, a method for realistically controlling the local lighting of an object through a coarse shadow, compatible with diffusion-based intrinsic image renderers without any additional training. Our results demonstrate that fine-grained control over local lighting can be achieved while attaining realistic compositions.
While our method can realistically generate relighting results from a shadow and despite showing that this can be done in multiple ways (shadow mapping, 2D generation, hand-drawn scribbles), the process of generating a shadow from scratch could prove challenging to some. A promising future direction is to enable end-to-end shadow control using parametric light models, such as point or sphere lights. Additionally, similar approaches could be explored to enable global lighting adjustments for entire images.
Acknowledgements. This research was supported by FRQNT scholarship 328810, NSERC grants RGPIN 2020-04799 and ALLRP 586543-23, Mitacs and Depix. Computing resources were provided by the Digital Research Alliance of Canada. The authors thank Zheng Zeng, Yannick Hold-Geoffroy and Justine Giroux for their help as well as all lab members for discussions and proofreading help.